OpenStudy (anonymous):

Jane poured 12 cups of water into a rectangular container measuring 12 inches by 8 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show Part C: What was the height of the water in the container? Show your work. (4 points)

OpenStudy (anonymous):

Only part c needed

OpenStudy (anonymous):

@dmart180

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@phi

OpenStudy (tkhunny):

See that "show your work" stuff? You should do that.

OpenStudy (anonymous):

I really don't understand how to do it

OpenStudy (anonymous):

@tkhunny I really just need an explanation

OpenStudy (tkhunny):

Well, then provide one. What's your plan? How did you do Part A? Where is Part B?

OpenStudy (tkhunny):

Total Volume = 12 * 8 * (9/2.54) CI Unfortunately, we are not told if 9 is the height. Is it? I guess we might be able to assume that it is, as it is given last. 12 cups = 12 * 144 CI = 12 * 8 * What? Work it out. Go without fear. You can't break the equation.

OpenStudy (anonymous):

5*5*5*3*...........................

OpenStudy (anonymous):

What ? @dieg

OpenStudy (anonymous):

sorrryyy